OpenStudy (anonymous):

A function is given below. Determine the average rate of change of the function between x = 1 and x = 1 + h. g(x) = -5/x

OpenStudy (anonymous):

basically, the average rate of change=total change/time the total change will be given by g(1+h)-g(1) and the time=1+h-1=h so, you should have g(1+h)=-5/(1+h), and g(1)=-5/1 the difference of these two divided by h shall be your answer